OpenStudy (anonymous):

Will fan and give medal if you help me answer these questions! 1. Your job pays $8 per hour. Write a variable expression for your pay in dollars working for H hours. What is your pay if you work 36 hours? a. 9h; $288 b. 8h; $288 c. 8h; $292 d. 8; $282 2. Evaluate the expression 84+3h, for h=2 a. 87 b. 90 c. 116 (Correct answer?) d. 174 3. Evaluate the expression A/4 for A+32 a. 9 b. 6 (Correct answer?) c. 8 d. 7 4. Evaluate the expression P+5-mn/2, for m=2, n=4 and p=1 a. 8 b. 4 c. 2 d. 6 And one more question

OpenStudy (anonymous):

1 is B... the answer is simple, multiply 8 by the number of hours you work, so 8h then substitute 36 for the h, so multiply 36(8) = 288

OpenStudy (anonymous):

5. Several people form groups with four people in each group. Write an expression for the number of groups formed by P people. How many groups are formed if there are 28 people a. 7p;4 b.28 (over) p;7 c. 4p; 28 d. p (over) 4; 7

OpenStudy (anonymous):

The rest are really easy so you should b able to figure them out... for 2 for example just solve 84 + 3(2) since h is 2 you just substitute it in

OpenStudy (anonymous):

I have math and reading dyslexia, and this is my first year of pre al, so this isn't easy for me.

OpenStudy (anonymous):

Ok well this isn't a quiz or test or anything right?

OpenStudy (anonymous):

no, it's just the ones I couldn't figure out from my book.

OpenStudy (anonymous):

ok well do you understand what I said for #1 and #2?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok so for #2 the one you thought was the right answer is not.... 3(2) is 6 so just add 6 to 84... 84 +6...so what do you get?

OpenStudy (anonymous):

90.

OpenStudy (anonymous):

yup! so there is your answer for 2... for #3 is that supposed to be a divided by 4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok lets come back to 3...we will do #4 right now...

OpenStudy (anonymous):

ok

OpenStudy (nirmalnema):

A/4 for A=32 32/4 = 8

OpenStudy (anonymous):

P+5-mn/2, for m=2, n=4 and p=1 1 + 5 - 2(4) 1 + 5 - 8 6 - 8 -2

OpenStudy (nirmalnema):

P+5-mn/2, for m=2, n=4 and p=1 1+5-(2*4)/2 6-4 =2

OpenStudy (anonymous):

oh wait 2 i meant 2

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (nirmalnema):

yup

OpenStudy (anonymous):

yes you should solve it like this P+5-mn/2, for m=2, n=4 and p=1 1 + 5 - 2(4)/2 1 + 5 - 8/2 1 + 5 - 4 6 - 4 2

OpenStudy (nirmalnema):

. p (over) 4; 7

OpenStudy (anonymous):

so you solve going right to left?

OpenStudy (anonymous):

Solve the multiplication/division part first....then do addition and subtraction

OpenStudy (anonymous):

Ok

OpenStudy (nirmalnema):

total no of people.=p people in each group=4 total groups = p/4

OpenStudy (anonymous):

have u learned PEMDAS?

OpenStudy (anonymous):

In like, 3rd grade. I don't remember tho

OpenStudy (nirmalnema):

if p=28 then total groups= 28/4 = 7

OpenStudy (anonymous):

Parentheses, exponents, multiplication/division, and addition/subtraction this is the order that you should solve equations

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

yea no problem :)

OpenStudy (anonymous):

For number 5 @nirmalnema had the correct answer

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

Thanks for the help!